Marvin’s Interiors issued 8-year bonds 2 years ago. The bonds have a face value of $1,800, a 7.0 percent, semiannual coupon, and a current market price of $1,389. What is the pre-tax cost of debt?

Multiple Choice:

  • 12.53 percent

  • 13.28 percent

  • 14.23 percent

  • 13.79 percent

The Up and Coming Corporation's common stock has a beta of 1.4. If the risk-free rate is 3.5 percent and the expected return on the market is 13 percent, what is the company's cost of equity capital?

Multiple Choice

  • 16.8%

  • 17.64%

  • 17.47%

  • 15.96%

  • 21.7%

Six months ago, you purchased 1,500 shares of ABC stock for $35.81 a share. You have received dividend payments equal to $0.70 a share. Today, you sold all of your shares for $38.95 a share. What is your total dollar return on this investment?

Multiple Choice

  • $4,710

  • $11,520

  • $1,050

  • $5,760

  • $6,729
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE07.09 PEx ENG. 16 01 2000 · FX E F G H J K L M N O D FOR YTM P 2 2 B C calculator strokes first of all set C/Y = P/Y = then N